OpenStudy (anonymous):

The square root of 2 is a rational number

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

Hint: All imperfect square roots are irrational numbers. Is the square root of 2 a perfect square?

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

So if you were to perform \(\sqrt{2}\) in your calculator, would you get a whole number?

OpenStudy (anonymous):

yes i got 1.41

OpenStudy (anonymous):

1.41 is not a whole number...

OpenStudy (anonymous):

ok so its false

OpenStudy (anonymous):

THat would mean that it is false because \(\sqrt{2}\) is not a rational number.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

of course the fact that you got a decimal by putting this into your calculator does not make it irrational

OpenStudy (anonymous):

for instance\[\sqrt{\frac19}\]will give you a decimal answer but is a rational number since it can be written as a ratio of two numbers\[\frac13\]